# RLTHF: Targeted Human Feedback for LLM Alignment

### v2

@@ 3.3 Reward Knowledge Transfer
RLTHF progressively converges toward the comprehensive human preference through iterative RM training and strategic human annotation investment. As shown in Figure  2(c) and  2(d), after five iterations, the reward distribution and accuracy curves closely align with the full-human annotation. Intermediate iteration curves can be found in Appendix B. The required number of iterations depends on the available human annotation and RM training budget. Notably, full-human alignment can sometimes be achieved before exhausting the annotation budget. In such cases, the samples selected for human annotation would largely lack distinct preference features, indicating that the model has effectively captured the human preference. Once desired alignment is achieved or the annotation budget is fully utilized, we proceed with fine-tuning the model for the downstream task. This can be done in two ways: 1) incorporating the final iteration RM into the PPO loop, or 2) labeling the whole dataset with the final RM and feeding the labeled dataset to a DPO pipeline.
@@ 4 Results
In this section, we present the results of our main experiments, conducted on two datasets: HH-RLHF (Bai et al., 2022a) and TL;DR (Völske et al., 2017). Specifically, we compare RLTHF against two baselines: (1) Random, where samples are randomly selected for human annotation (matching the number of human-annotated samples in RLTHF), with the rest relying on AI feedback, and (2) Human, where all samples are annotated by humans. A detailed description of our experimental setup is provided in Appendix C.
@@ 4.1.1 Overall Alignment Improvement
Figure 3: Overall preference accuracy improvement of RLTHF on test data in an iterative manner. Here, we experiment with different sizes of down-sampled training data shards for RLTHF. We find 6% (HH-RLHF) to 7% (TL;DR) total human annotations on 1/4 shard yielding the optimal alignment for downstream tasks.
We here use GPT-4o for the initial alignment and evaluate RLTHF’s iterative alignment improvements by measuring the preference accuracy of RMs trained with varying proportions of human annotations relative to the full dataset. We employ RLTHF on both the complete dataset and multiple down-sampled shards as described in § 3.2.5. For a given shard, we run RLTHF in an iterative manner infusing targeted human annotations in each iteration. We evaluate the trained RMs on a separate test dataset, ensuring a clear distinction from the training data, and report their preference accuracy in Figure 3.
In Figure 3, each data point for a shard corresponds to an iteration of RLTHF. The results show a consistent improvement in test preference accuracy across iterations, with significant early gains that gradually diminish as accuracy approaches the upper bound. Additionally, down-sampling enhances the efficiency of human annotations: RLTHF running on 1/2 and 1/4 shards outperforms its full-dataset counterpart when using the same number of human annotations. However, excessive down-sampling (e.g., 1/8 shard) may limit the achievable accuracy due to reduced data richness. For downstream task fine-tuning, we identify 1/4 shard as the optimal choice. Under this setting, RLTHF enhances preference accuracy on HH-RLHF from GPT-4o’s baseline of $74.7\%$ to $89.6\%$ with only $6\%$ human annotations, and on TL;DR, from $78.8\%$ to $88.0\%$ with just $7\%$ human annotations. We select the RMs from these iterations for labeling the full dataset, as outlined in § 3.2.5. Evaluating the test accuracy of RMs trained on the fully labeled dataset, we observe further improvements, with accuracy reaching $91.8\%$ for HH-RLHF and $89.6\%$ for TL;DR.
@@ 4.1.2 Comparison against the Baselines
We begin by using two different LLMs—GPT-4o and GPT-4o mini—for the initial AI labeling. We then employ two separate RLTHF pipelines, RLTHF (4o) and RLTHF (4o mini), to improve alignment. To evaluate their effectiveness, we compare these pipelines against two baselines: (1) AI-only labeling (GPT-4o and GPT-4o mini), (2) Random human annotation, and (3) fully Human-annotated datasets (baseline setup details are provided in Appendix C.3).
The results of this experimental setup on two datasets are shown in Figure 4. RLTHF (4o) consistently outperforms Random (4o), as random human annotation proves ineffective in correcting AI mislabeling, resulting in only marginal improvements in test accuracy. Of particular interest is the ”Return on Investment (ROI)”—measured as the increase in test accuracy per unit of human annotation. With just $6\%$ human annotation, RLTHF (4o) achieves a 15.9×15.9$\times$15.9 × and 5.3×5.3$\times$5.3 × higher ROI compared to Random (4o) on HH-RLHF and TL;DR, respectively.
Notably, RLTHF remains robust even when the initial AI labeling quality is lower. While GPT-4o mini starts with an accuracy gap of $2.6\%$ and $5.6\%$ compared to GPT-4o, this gap shrinks to just $0.4\%$ and $-0.2\%$ after incorporating $10\%$ human annotation on HH-RLHF and TL;DR, respectively. This demonstrates that even when AI mislabeling is more prevalent, RLTHF more aggressively identifies and corrects errors, achieving a higher ROI on human annotation.
Figure 4: RLTHF effectively bridges the prefernce accuracy gap between LLM-generated (GPT-4o/GPT-4o-mini) labels and fully human-annotated data, regardless of the initial labeling accuracy of the LLM. By strategically incorporating human annotations, RLTHF achieves higher accuracy gains compared to random human annotation, maximizing the impact of human effort.
@@ 4.1.3 Effects of Hyperparameters
(a) Amplification Ratio (HH-RLHF)
(b) Back-off Ratio (HH-RLHF)
(c) Annotation Batch Size (HH-RLHF)
@@
Figure 5: Effects of RLTHF hyperparameters on HH-RLHF (top row) and TL;DR (bottom row). A larger amplification ratio ($\alpha$) and back-off ratio ($\beta$) are beneficial in the initial iterations but should be gradually reduced as human annotations accumulate. The human annotation budget should be distributed across multiple batches, though the impact of finer granularity remains unclear.
Amplification Ratio. To investigate how the human amplification ratio $\alpha$ contributes to RLTHF, we fix the back-off ratio $\beta$ at 60% and conduct a controlled study on different amplification ratios.
The results for each dataset are shown in Figures 5(a) and 5(d). We observe that both no amplification ($\alpha=1$) and excessive amplification ($\alpha=8$) of human annotations lead to suboptimal RM improvements. Specifically, lower amplification results in smaller improvements in the initial iterations, while in later iterations, this trend reverses. This is expected, as no or low amplification weakens the impact of human annotations, particularly in the early iterations when the total number of annotations remains low, while over-amplification skews the training data distribution and increases the risk of overfitting, especially in later iterations when the base number of annotations is already large.
Although we use a fixed amplification ratio (Appendix C.4), we recommend starting with a higher amplification ratio and gradually reducing it as the number of accumulated annotations grows to achieve the best results.
Itr # HH-RLHF TL;DR
@@
Table 1: Ablation study on HH-RLHF and TL;DR. All three factors, human annotation, amplification of human annotations, and the back-off mechanism, play a crucial role in RLTHF’s effectiveness.
Back-off Ratio. To investigate how the back-off ratio $\beta$ contributes to RLTHF, we conduct a controlled study while keeping all other hyperparameters at their default values.
The results for each dataset are shown in Figures 5(b) and 5(e). We observe a clear trend where a larger $\beta$ leads to higher improvements in the initial iteration, but its gains diminish in later iterations. In contrast, a smaller $\beta$ results in slower improvements early on but accelerates in later iterations, reflecting a shift in the trade-off between data richness and quality. At the beginning, the data is relatively unsanitized, making data quality the limiting factor, so a higher back-off ratio is beneficial due to its stronger sanitization effect. However, as accuracy improves over time, even a smaller back-off ratio can produce sufficiently clean data, allowing its greater data richness and diversity to become the dominant advantage.
Similar to our settings (Appendix C.4), we recommend starting with a higher back-off ratio and gradually reducing it as the data becomes more sanitized.
Annotation Batch Size. To investigate how the number of annotated samples per iteration affects RLTHF, and more fundamentally, whether RLTHF benefits from an iterative approach compared to a one-shot annotate-all strategy, we conduct a controlled study while keeping the amplification ratio $\alpha$ fixed at 4 and the back-off ratio $\beta$ at 60%.
The results for each dataset are shown in Figures 5(c) and 5(f) (note that the annotation percentages are relative to the shard rather than the full dataset). We observe that the iterative approach achieves an improvement of up to $4.2\%$ by Itr-4 compared to the one-shot annotation strategy. This suggests that in each iteration, the RM not only learns from the annotated samples but also generalizes to a broader set of similar samples.
However, in our preliminary study, we did not observe a significant benefit when splitting the annotation budget into finer batches (e.g., 4% per iteration vs. 8% per iteration). Given the additional GPU time and human feedback round trips required for smaller batches, we recommend splitting the annotation budget into two batches or using a similarly coarse granularity to balance efficiency and performance.
@@ 4.1.4 Ablation Study
We conduct an ablation study to evaluate the necessity of each component in RLTHF and understand their contributions. All experiments in this study use a 4×4$\times$4 × amplification ratio, a $60\%$ back-off ratio, and $4\%$ human annotations relative to a 1/4 shard per iteration. Our primary goal is to answer two key questions.
Does self-improvement alone work? To assess whether RLTHF can function effectively without human annotations, we set the annotation batch size to $0\%$, reducing RLTHF to a pure self-improvement scheme. The results in Table 1 indicate that self-improvement alone has inherent limitations and struggles to surpass GPT’s baseline preference accuracy. Further evaluation on downstream tasks also confirms that this GPT-level preference accuracy is insufficient for aligning with human preferences. These findings suggest that self-improvement alone is inadequate for RLTHF’s intended use case.
Are amplification and back-off necessary? To examine the importance of RLTHF’s human amplification and sanitization back-off mechanisms, we disable both simultaneously (since the individual ablation of amplification was already tested in § 4.1.3 with $\alpha=1$). The results in Table 1 show that RLTHF achieves only marginal improvement without these mechanisms, even with accumulating annotations. This result highlights the strong contributions of both amplification and back-off. Without sanitization back-off, incorrect labels persist in the training data, making it harder to identify valuable samples and further reducing data cleanliness. Without human amplification, an even greater number of correct samples is required to overturn incorrect ones, exacerbating the issue.
@@ 4.2 Downstream Tasks
To assess RLTHF’s effectiveness on final downstream tasks, we conduct DPO training on RLTHF-prepared data and compare it against other baselines w.r.t. an SFT model. For RLTHF, we leverage the RM trained on a 1/4 shard, refined with targeted $6\%$ human annotation for HH-RLHF and $7\%$ for TL;DR, to regenerate preferences for the full dataset. We then train an LLM with DPO on this refined dataset. For the Random baseline, we introduce the same amount of randomly selected human annotations on top of AI labels into the DPO training dataset. Performance is evaluated using the pairwise win rate against SFT model outputs, calculated with AlpacaEval (Li et al., 2023b) running Claude 3.5 Sonnet (Anthropic, 2024) as a judge on test sets from HH-RLHF and TL;DR datasets (see Appendix D for details).
Dataset HH-RLHF TL;DR
AI Labeler GPT-4o GPT-4o GPT-4o mini
AI-labeled 49.2 59.2 56.4
Random 52.5 59.8 57.5
RLTHF 58.1 62.3 62.4
@@
Table 2: Win rate against SFT (%). RLTHF consistently outperforms the AI-labeled, Random, and Human baselines across both datasets. Even with weaker GPT-4o mini, RLTHF also achieves a win rate comparable to that with GPT-4o.
@@ A.3 User Prompt for TL;DR
Appendix B Iterative Alignment Improvement

### v3

@@ 3.3 Reward Knowledge Transfer
RLTHF progressively converges toward the oracle human preference through iterative RM training and strategic human annotation investment. As shown in Figure  2(c) and  2(d), after five iterations, the reward distribution and accuracy curves closely align with the full-human annotation. Find the intermediate iteration curves in Appendix C. The required number of iterations depends on the available human annotation and RM training budget. Notably, full-human alignment can sometimes be achieved before exhausting the annotation budget. In such cases, the samples selected for human annotation would largely lack distinct preference features, indicating that the model has effectively captured the human preference. Once desired alignment is achieved or the annotation budget is fully utilized, we proceed with fine-tuning an LLM for the downstream task. This can be done in two ways: 1) incorporating the final iteration RM into the PPO loop, or 2) labeling the whole dataset with the final RM and feeding the labeled dataset to a DPO pipeline.
@@ 4 Results
In this section, we present the results of our main experiments, conducted on two datasets: HH-RLHF (Bai et al., 2022a) and TL;DR (Völske et al., 2017). Specifically, we compare RLTHF against three baselines: (1) AI-only labeling, where samples are fully labeled by LLMs, (2) Random human annotation, where samples are randomly selected for human annotation (matching the number of human-annotated samples in RLTHF), with the rest relying on AI feedback, and (2) Human, where all samples are annotated by humans. A detailed description of our experimental setup is provided in Appendix D.
@@ 4.1.1 Overall Alignment Improvement
Figure 3: Overall preference accuracy improvement of RLTHF on test data in an iterative manner. Here, we experiment with different sizes of down-sampled training data shards for RLTHF. We find 6% ($H$-RLHF) to 7% (TL;DR) total human annotations on 1/4 shard yielding the optimal alignment for downstream tasks.
We here use GPT-4o for the initial alignment and evaluate RLTHF’s iterative alignment improvements by measuring the preference accuracy of RMs trained with varying proportions of human annotations relative to the full dataset. We employ RLTHF on both the complete dataset and multiple down-sampled shards as described in § 3.2.5. For a given shard, we run RLTHF in an iterative manner infusing targeted human annotations in each iteration. We evaluate the trained RMs on a separate test dataset and report their preference accuracy in Figure 3.
In Figure 3, each data point for a shard corresponds to an iteration of RLTHF. The results show a consistent improvement in test preference accuracy across iterations, with significant early gains that gradually diminish as accuracy approaches the upper bound. Additionally, down-sampling enhances the efficiency of human annotations: RLTHF running on 1/2 and 1/4 shards outperforms its full-dataset counterpart when using the same number of human annotations. However, excessive down-sampling (e.g., 1/8 shard) may limit the achievable accuracy due to reduced data richness. For downstream task fine-tuning, we identify 1/4 shard as the optimal choice. Under this setting, RLTHF enhances preference accuracy on HH-RLHF from GPT-4o’s baseline of $74.7\%$ to $89.6\%$ with only $6\%$ human annotations, and on TL;DR, from $78.8\%$ to $88.0\%$ with just $7\%$ human annotations. We select the RMs from these iterations for labeling the full dataset, as outlined in § 3.2.5. Evaluating the RMs trained on the fully human-labeled dataset, we observe the accuracy to be $91.8\%$ for HH-RLHF and $89.6\%$ for TL;DR.
@@ 4.1.2 Comparison against the Baselines
We begin by using two different LLMs – GPT-4o and GPT-4o mini – for the initial AI labeling. We then employ two separate RLTHF pipelines, RLTHF (4o) and RLTHF (4o mini), to improve alignment. To evaluate their effectiveness, we compare these pipelines against three aforementioned baselines (details in Appendix D.3).
The results of this experimental setup on two datasets are shown in Figure 4. RLTHF (4o) consistently outperforms Random (4o), as random human annotation proves ineffective in correcting AI mislabeling, resulting in only marginal improvements in test accuracy. Of particular interest is the “Return on Investment (ROI)”, which is measured as the increase in test accuracy per unit of human annotation. With just $6\%$ human annotation, RLTHF (4o) achieves a $15.9\times$ and $5.3\times$ higher ROI compared to Random (4o) on HH-RLHF and TL;DR, respectively.
Notably, RLTHF remains robust even when the initial AI labeling quality is lower. While GPT-4o mini starts with an accuracy gap of $2.6\%$ and $5.6\%$ compared to GPT-4o, this gap shrinks to just $0.4\%$ and $0.2\%$ after incorporating $10\%$ human annotation on HH-RLHF and TL;DR, respectively. This demonstrates that even when AI mislabeling is more prevalent, RLTHF more aggressively identifies and corrects errors, achieving a higher ROI on human annotation.
Figure 4: RLTHF effectively bridges the prefernce accuracy gap between LLM-generated (GPT-4o/GPT-4o-mini) labels and fully human-annotated data, regardless of the initial labeling accuracy of the LLMs. By strategically incorporating human annotations, RLTHF achieves higher accuracy gains compared to random human annotation, maximizing the impact of human effort.
@@ 4.1.3 Effects of Hyperparameters
(a) Amplification Ratio ($H$-RLHF)
(b) Back-off Ratio ($H$-RLHF)
(c) Annotation Batch Size ($H$-RLHF)
@@
Figure 5: Effects of RLTHF hyperparameters on HH-RLHF (top row) and TL;DR (bottom row). A larger amplification ratio ($\alpha$) and back-off ratio ($\beta$) are beneficial in the initial iterations but should be gradually reduced as human annotations accumulate and dataset gets sanitized. The human annotation budget should be distributed across multiple batches, though the impact of finer granularity remains unclear.
Amplification Ratio. To investigate how the amplification ratio (discussed in Section 3.2.5) $\alpha$ contributes to RLTHF, we fix the back-off ratio $\beta$ at $60\%$ and conduct a controlled study on different amplification ratios. The results for each dataset are shown in Figures 5(a) and 5(d). We observe that both no amplification ($\alpha=1$) and excessive amplification ($\alpha=8$) of human annotations lead to suboptimal RM improvements. Specifically, lower amplification results in smaller improvements in the initial iterations, while in later iterations, this trend reverses. This is expected, as no or low amplification weakens the impact of human annotations, particularly in the early iterations when the total number of annotations remains low, while over-amplification skews the training data distribution and increases the risk of overfitting, especially in later iterations when the base number of annotations is already large. For the best results, we start with a higher value of $\alpha$ and gradually reduce it (Appendix D.4).
Itr # $H$-RLHF TL;DR
@@
Table 1: Ablation study of RLTHF on HH-RLHF and TL;DR. All three factors – human annotation, amplification of human annotations, and the back-off mechanism – play a crucial role in RLTHF’s effectiveness.
Back-off Ratio. To investigate the role of the back-off ratio $\beta$ RLTHF, we conduct a controlled study where all other hyperparameters are held at their default values. The results across datasets are presented in Figures 5(b) and 5(e). We observe a consistent pattern: a larger $\beta$ yields greater improvement during the initial iteration, but its benefit diminishes in later iterations. Conversely, a smaller $\beta$ leads to slower initial progress but accelerates in later stages. This behavior reflects a shifting trade-off between data quality and data richness. Early in training, when the dataset is relatively unsanitized, quality is the primary bottleneck—making a higher back-off ratio advantageous due to its stronger filtering effect. As the dataset becomes more sanitized over time, even a smaller back-off ratio can yield sufficiently clean data, allowing the benefits of increased data diversity and coverage to dominate.
In line with our configuration settings (Appendix D.4), we recommend starting with a high back-off ratio and progressively reducing it as the dataset becomes more sanitized.
Annotation Batch Size. To evaluate how the number of annotated samples per iteration impacts the effectiveness of RLTHF – and more broadly, to assess whether RLTHF benefits from an iterative strategy over a one-shot annotate-all approach – we conduct a controlled study, fixing the amplification ratio $\alpha$ at 4 and the back-off ratio $\beta$ at 60%.
The results, shown in Figures 5(c) and 5(f) (with annotation percentages computed relative to the shard rather than the full dataset), indicate that the iterative approach yields up to a $4.2\%$ improvement by Itr-4 compared to the one-shot annotation strategy. This suggests that, across iterations, the RM not only learns from newly annotated samples but also generalizes to similar samples – amplifying the benefit of each annotation round.
However, our preliminary experiments show that further subdividing the annotation budget into smaller batches (e.g., 4% vs. 8% per iteration) does not provide a meaningful advantage. Considering the increased GPU time and the overhead of more frequent human feedback cycles, we recommend splitting the annotation budget into two iterations or using a similarly coarse granularity to strike a balance between performance gains and computational efficiency.
@@ 4.1.4 Ablation Study
We conduct an ablation study to assess the necessity of two core mechanisms in RLTHF and understand their individual contributions. All experiments in this study are performed with $\alpha=4$, $\beta=60\%$, and $4\%$ human annotations per iteration, relative to a 1/4 shard of the dataset. The primary objective of this study is to address two key questions.
Does self-improvement alone work? To evaluate whether RLTHF can operate effectively without human annotations, we set the annotation batch size to $0\%$, effectively reducing RLTHF to a purely self-improving framework, as proposed in recent studies (Huang et al., 2024; Yu et al., 2025). The results, shown in Table 1, reveal that self-improvement alone faces inherent limitations and fails to exceed the baseline preference accuracy achieved by the AI (GPT in this case). Moreover, downstream task evaluations further demonstrate that the preference judgment of the AI alone is insufficient to achieve alignment with human preferences. These findings indicate that self-improvement based solely on AI’s preference is inadequate for oracle human alignment.
Are amplification and back-off necessary? To assess the importance of RLTHF’s human annotation amplification and sanitization back-off mechanisms, we ablate both mechanisms simultaneously – given that the effect of removing amplification alone was already examined in § 4.1.3 using $\alpha=1$. As shown in Table 1, RLTHF achieves only marginal improvement under this configuration, even as human annotations accumulate over iterations. Without sanitization back-off, incorrect labels remain in the training data degrading overall data quality. Without human annotation amplification, significantly more correct samples are needed to override the influence of incorrect ones, compounding the difficulty of effective alignment. Together, these results highlight that both mechanisms are essential for driving meaningful improvements in RLTHF.
@@ 4.2 Downstream Tasks
To evaluate RLTHF’s effectiveness on final downstream tasks, we perform DPO training using data prepared by RLTHF and compare it against three baselines. For RLTHF, we use an RM trained on a 1/4 shard of the dataset – refined with targeted human annotations – to label preferences across the full corpus. Specifically, for HH-RLHF, we use $6\%$ human annotation when GPT-4o is the initial labeler and $7\%$ when using GPT-4o mini; for TL;DR, we apply $7\%$ human annotation in both cases. We then train an LLM, Qwen2.5-3B, with DPO using the RLTHF-refined dataset. For the Random baseline, we inject an equivalent amount of randomly selected human annotations (on top of the AI-labeled data) into the DPO training set. Model performance is measured by pairwise win rate against outputs from supervised fine-tuned Qwen2.5-3B, using AlpacaEval (Li et al., 2023b) with Claude 3.5 Sonnet (Anthropic, 2024) as the evaluation judge. Evaluations are conducted on held-out test sets from both datasets (details in Appendix E).
Dataset $H$-RLHF TL;DR
AI Labeler 4o 4o-mini 4o 4o-mini
AI-labeled 49.2 45.1 59.2 56.4
Random 52.5 46.3 59.8 57.5
RLTHF 58.1 56.1 62.3 62.4
@@
Table 2: Win rate against SFT (%). RLTHF outperforms the AI-labeled, Random, and Human baselines across both datasets. Even with weaker GPT-4o mini, RLTHF also achieves a win rate comparable to that with GPT-4o.
@@
4.3 Cost Analysis
RLTHF introduces two types of additional costs: (1) the cost of LLM annotation during initial alignment, and (2) the cost of iterative RM training. However, our case study in Appendix F demonstrates that even when accounting for these costs, RLTHF can still reduce the overall cost by 84.0–86.0%, based on a conservative estimate.
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.3 User Prompt for TL;DR
Appendix B Pseudocode for the Full RLTHF Procedure
In Algorithm 1, we present the full procedure of RLTHF in pseudocode.
Algorithm 1 Pseudocode for the Full RLTHF Procedure
[…]
Output: Final aligned labels $\mathcal{L}$ or reward model $RM_{\text{final}}$
 // Step 1: Initial Alignment on a Random Shard
$S_{0}\leftarrow D.\texttt{randomSubset}()$
$\mathcal{L}_{0}\leftarrow S_{0}.\texttt{labelWith}(LLM)$
$RM_{0}\leftarrow\texttt{TrainRewardModel}(\mathcal{L}_{0})$
 // Step 2: Iterative Alignment Improvement
for $i=0$ to M1M-1 do
  $\Delta_{i}\leftarrow\texttt{EmptyList}()$
  for all $(x,y_{c},y_{r})\in\mathcal{L}_{i}$ do
   $\delta\leftarrow RM_{i}(x,y_{c})-RM_{i}(x,y_{r})$
   $\Delta_{i}.\texttt{append}(\langle x,y_{c},y_{r},\delta\rangle)$
  end for
  $\mathcal{S}_{\text{sorted}}\leftarrow\Delta_{i}.\texttt{sortBy}(\delta)$
  $\mathcal{C}_{i}\leftarrow\mathcal{S}_{\text{sorted}}.\texttt{findCorrectSamples}()$
  $\mathcal{H}_{i}\leftarrow\mathcal{S}_{\text{sorted}}.\texttt{findHardSamples}().\texttt{labelWith}(Human)$
  $\mathcal{R}_{i}\leftarrow\mathcal{S}_{\text{sorted}}.\texttt{findIncorrectSamples}().\texttt{flipLabels}()$
  $\mathcal{L}_{i+1}\leftarrow\mathcal{C}_{i}\cup\mathcal{R}_{i}\cup\mathcal{H}_{i}$
  $RM_{i+1}\leftarrow\texttt{TrainRewardModel}(\mathcal{L}_{i+1})$
end for
 // Step 3: Extend to Full Dataset
$\mathcal{L}\leftarrow D.\texttt{labelWith}(RM_{M})$
$RM_{\text{final}}\leftarrow\texttt{TrainRewardModel}(\mathcal{L})$
 // Step 4: Knowledge Transfer
Option 1: Use $\mathcal{L}$ to directly align a downstream model (e.g., DPO)
Option 2: Integrate $RM_{\text{final}}$ into RL-based optimization (e.g., PPO)
Appendix C Iterative Alignment Improvement